Opened 9 years ago

Closed 8 years ago

Last modified 12 months ago

#284 closed flight processing (fixed)

ET07/01, flight day 300a/2009, Afar north

Reported by: benj Owned by: knpa
Priority: alpha 5 Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by mark1)

Data location: ~arsf/arsf_data/2009/flight_data/ethiopia/ET07_01-2009_300a_Afar_north

Data arrived from ARSF via SATA disk Y on 4th Nov. 2009.

Scientific objective: /users/rsg/arsf/arsf_data/2008/ARSF_Applications/Ethiopia/ET07_01\ -\ Dr\ Wright.pdf

Priority: alpha-5

PI: Tim Wright

2009 Ethiopia campaign, reflown from 2007

Sensors:

  • Eagle (26/02/2010)
  • Hawk (26/02/2010)
  • Leica LIDAR (26/02/2010)
  • Digital Photography (26/02/2010)
  • TABI (to be processed by University of Athens)

Change History (25)

comment:1 Changed 9 years ago by benj

  • Description modified (diff)

The aircraft had an engine power-loss on landing, which meant the equipment got powered-down prematurely. This resulted in some of the Lidar data being corrupted. The affected data was re-flown on day 301.

Also, when I checked it in Ethiopia there was some corruption of the Applanix data (something bad in dcalm.005 I think it was). However, the Applanix data from flight 300b seems to also contain the data from 300a, so if there is a problem with the Applanix data from this flight then use the data from 300b.

comment:2 Changed 9 years ago by benj

  • Description modified (diff)

Raw GPS base station data is in ~arsf/arsf_data/2009/flight_data/ethiopia/basestations.

comment:3 Changed 9 years ago by mggr

  • Description modified (diff)

Raw TABI data acquired, put in archive and offered to Athens for download and processing.

comment:4 Changed 9 years ago by emca

  • Owner set to emca
  • Status changed from new to assigned

comment:5 Changed 9 years ago by mark1

Photograph delivery directory created. Photos 1-113 tagged by IPAS SOL file, 114-330 tagged using log file (since IPAS or LIDAR failed mid flight).

comment:6 Changed 9 years ago by mark1

Photographs checked and ready to deliver.

comment:7 Changed 9 years ago by emca

LIDAR processing is complete and ready for delivery checking at
~airborne/workspace/ET07_01-2009_300a_Afar_north/delivery/20100128

comment:8 Changed 9 years ago by emca

Eagle and Hawk SCT offsets per line

Line Eagle Hawk
1-0.01 -0.07
2+0.03 +0.01
3-0.03 -0.02
4+0.04 0.00

comment:9 Changed 9 years ago by emca

Hyperspectral processing complete and ready for delivery checking at
~airborne/workspace/ET07_01-2009_300a_Afar_north/delivery/20100215

comment:10 Changed 9 years ago by mark1

LIDAR delivery checked and all looks OK.

comment:11 Changed 9 years ago by mark1

Hyperspectral data delivery checked. Example commands tested with dem and work OK, all level 1 files checked and look good. Minor edits made to read-me file.

comment:12 Changed 9 years ago by mark1

  • Description modified (diff)

Eagle, Hawk and LIDAR data delivered to PI on 26/02/2010

comment:13 Changed 9 years ago by mark1

  • Description modified (diff)

Digital photography also delivered on 26/02/2010

comment:14 Changed 9 years ago by benj

Processed TABI data was received from Pol back in January but we weren't happy with the accuracy. This data is in the workspace, it has not been dispatched to the PI but is the only processed TABI data we currently have for this project.

comment:15 Changed 9 years ago by emca

  • Component changed from Processing: general to Archiving

All data has been rsynced to
~arsf/arsf_data/2009/flight_data/ethiopia/ET07_01-2009_300a_Afar_north

Ready for archiving

comment:16 Changed 8 years ago by knpa

  • Owner changed from emca to knpa

Beginning archiving.

comment:17 Changed 8 years ago by knpa

Some ipas files seem to be missing (intermediate files, usually in gps folder).

comment:18 Changed 8 years ago by knpa

Navigation data is a bit confusing.

For applanix, it looks like it was processed three times. Have separated proc data which appears to be related to previous runs into applanix/proc/sbet_old. However, data in applanix/extract and applanix/gps seem to relate to older runs only. Have left these as they are.

For ipas, the extract folder is odd in that all files have a modification time of June, four months after delivery. As mentioned before, the gps data is missing.

comment:19 Changed 8 years ago by knpa

Only 1 lidar line included in delivery.

From Readme:
"Please note that due to a technical fault, this delivery only includes LIDAR
data for line Nth3. The remaining lines (Nth5, Nth4 and Nth2) were re-flown
on day 301"

comment:20 Changed 8 years ago by knpa

hyperspectral delivery is missing overview screenshots (i.e. eagle/hawk mosaics).

comment:21 Changed 8 years ago by knpa

Finished preparing for archiving.

Larsen is currently full, so cannot create tarball yet.

comment:22 Changed 8 years ago by knpa

Sent to NEODC, and backed up by them.

Have moved out of repository, to archive disk.

comment:23 Changed 8 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleted workspace copy.

Closing ticket.

comment:24 Changed 2 years ago by lah

Tidying

There is no reprocessed hyperspectral delivery, but there are reprocessed files, so I've created a new delivery out of these. There was a readme already generated so I've copied this into the directory. This will need a delivery check before archiving. I have not deleted any of the hyperspectral processing files yet - will wait until after the delivery is checked.

comment:25 Changed 12 months ago by asm

Archiving

Reprocessed hyperspectral archived under: http://data.ceda.ac.uk/neodc/arsf/2009/ET07_01/ET07_01-2009_300a_Afar_north/ET07_01-300a-hyperspectral-20161028/

It has not replaced the old delivery but all data is present so will mark the project as fixed.

Note: See TracTickets for help on using tickets.